OpenStudy (anonymous):

evaluate 7xfor x= 2/3 a. 10 and 1/2 b. 2/21 c. 3/14 d. 4 and 2/3? I think its d. evaluate 5/3 divided by 4/n for n= 3 a. 1 and 1/4 b. 2 and 2/9 c. 4/5 d. 9/20 I think a?

OpenStudy (anonymous):

what do you think 7x means? just tell me your answer in english, no need for math.

OpenStudy (anonymous):

yes D

OpenStudy (anonymous):

I think its four and two thirds

OpenStudy (anonymous):

yes, thats right.

OpenStudy (anonymous):

Thanks, and is the second one a??? one and one fourth

OpenStudy (anonymous):

(5/3)/(4/3) = (5/3) * (3/4) = (5/4) = 1.25 which is 1 and 1/4

OpenStudy (anonymous):

YES!!!!! Thanks for helping guys, I wasnt sure of myself as usual o_o

OpenStudy (anonymous):

I'm not sure number 2 question (5/3)/(4/3)?

OpenStudy (anonymous):

yes. I expanded it above.

OpenStudy (anonymous):

its 5/3 but you 4/n but we need to find the N for that 4/n which is 1 and 1/4

OpenStudy (anonymous):

oh i see what you meant. it is given that n = 3 so (5/3)/(4/n) = (5/3)/(4/3)

OpenStudy (anonymous):

Yeppie :)